AOL-TWX merger Question

On Jan 10 2000 AOL announced acquisition of Time Warner (TWX), where 1.5 shares of AOL were exchanged for 1 share of TWX. Table shows pre-merger values:

Share price number of shares market cap (pre-merger)
AOL $72.88 2.6 billion $189.5B
TWX $64.75 1.4 billion $90B

a) what was the % ownership split of Shareholders after transaction?

b) Assume there are no synergies between the firms. What was the % premium that TWX Shareholders received? How much would AOL have to pay TWX in cash such that shareholders of TWX would be indifferent between share or cash acquisition?

c) On July 18 2002, the stock price (of combined company) was $12.75. How much value was created since the acquisition of TWX? How is value creation or destruction shared between the (former) shareholders of AOL and the (former) sharesholders of TWX? What if alternatively AOL had bought TWX using cash?

d) looking back, would the (Former) sharesholders of TWX have preferred a stock or cash acquisition in (c)?
Answer #1

1) Post-transaction there would be a total of 4.7 billion shares (=1.5*1.4 billion + 2.6 billion), and the former shareholders of AOL owned 55.32% (= 2.6/4.7) of the company, and the former shareholders of TWX owned 44.68% (=2.1/4.7) of the company

2) Premium = 1.5 * PAT - PTWX.

PAT = VAT / (Total number of shares) = $280.2/4.7 = $59.62

Premium =1.5 * $59.62- $64.75 = $24.68.

Premium (in %) = $24.68 / $64.75 = 38%

Cash value = $59.62 * 2.1 Billion = $125.2 Billion

Thus the shareholders of AOL are indifferent to buy TWX cash at $125.2 Billion or to perform a share acquisition with a conversion of share of 1 TWX equals 1.5 AOL.

3) On July 18, 2002 the company's value was $59.9 billion (=$12.75 * 4.7 billion). Since January 10, 2000 the value was destroyed by 220.3 billion (= $280.2 billion - $59.9 billion). The former AOL shareholders faced a loss for 55.32% of the loss i.e. $121.87 billion, and the former TWX shareholders suffered a loss for 44.68% i.e. $98.43 billion. When the acquisition was performed by cash, than AOL shareholders would have suffered 100% of the loss (in an acquisition by cash, the target company's shareholder's do not take any part of future profits/losses because they exit the organisation).

4) The TWX shareholders ex-post would have preferred a cash acquisition